# Generalized quantum Zernike Hamiltonians: Polynomial Higgs-type algebras and algebraic derivation of the spectrum

### v1

@@ 2 Generalized quantum Zernike Hamiltonians
  • Introduce number and ladder (raising/lowering) operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ from $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$ via a nonlinear change of basis (which depends on the chosen $N$ through the parameters $\gamma_{k}$), such that
    $\bigl{[}\hat{\mathcal{K}},\hat{\mathcal{K}}_{\pm}\bigr{]}=\pm\hat{\mathcal{K}}_{\pm},\qquad\bigl{[}\hat{\mathcal{K}}_{-},\hat{\mathcal{K}}_{+}\bigr{]}=\Phi\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}+{\rm Id}\bigr{)}-\Phi\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}\bigr{)},$ (2.9)
@@
    $\hat{\mathcal{K}}_{+}\hat{\mathcal{K}}_{-}\equiv\Phi\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}\bigr{)}=\Phi_{1}\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}\bigr{)}\Phi_{2}\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}\bigr{)}.$ (2.10)
@@
  • Finally, using the eigenvalues $E$ of ^$\hat{\mathcal{H}}_{N}$ and ${B}$ of ${\hat{B}}$, we consider a finite-dimensional representation of (2.12), that is,
@@
    for a natural number ${n}\in\{1,2,\dots\}$. This set of algebraic equations provides us with constraints on the spectrum and the representation-dependent constant $u$. Hence the solutions take the form $u=u({n})$ and $E=E({n})$, thus leading to $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}\in\{1,\dots,{n}\}$. If $\Phi(B,{n})>0$, the finite-dimensional representation is also unitary. The parameter ${n}$, which characterizes the dimension of the subspace, i.e. of dimension ${n}+1$, determines the principal quantum number $\mathfrak{n}$.
In the following sections 35, we work out the quadratic, cubic and quartic generalized quantum Zernike Hamiltonians ^$\hat{\mathcal{H}}_{N}$ (2.3), corresponding to the values $N=2,3,4$, respectively. For each case, we obtain the quantum symmetries (2.7), the polynomial algebras (2.8) and (2.9), together with the deformed oscillator algebra (2.12). From the latter, the solutions for the possible spectra determined by the equations (2.14) are deduced and analyzed. Furthermore, such new results are also interpreted as superintegrable perturbations of the isotropic oscillator on the two-dimensional sphere, hyperbolic and Euclidean spaces.
While we have not yet fully derived the general expressions for the deformed oscillator algebra (2.12) along with the corresponding spectra for arbitrary $N$, we present two conjectures in section 6 that address these problems. Additionally, we provide a brief overview of the case $N=5$ to further verify the correctness of the results. These conjectures serve as a foundation for future work to establish their proofs.
@@ 3 The proper (quadratic) quantum Zernike Hamiltonian
Let us consider the quadratic case of ^$\hat{\mathcal{H}}_{N}$ (2.3) with $N=2$ and arbitrary parameters $\gamma_{1}$ and $\gamma_{2}$. The second-order in the momenta quantum symmetries (2.6) commuting with ^$\hat{\mathcal{H}}_{2}$ (2.7) are found to be
$\begin{split}{\hat{I}}_{2}&={\hat{p}}_{2}^{2}+\gamma_{1}{\hat{q}}_{2}{\hat{p}}_{2}-\gamma_{2}\bigl{(}{\hat{q}}_{2}^{2}({\hat{p}}_{1}^{2}-{\hat{p}}_{2}^{2})-2{\hat{q}}_{1}{\hat{q}}_{2}{\hat{p}}_{1}{\hat{p}}_{2}+{\rm i}{\hat{q}}_{1}{\hat{p}}_{1}+{\rm i}{\hat{q}}_{2}{\hat{p}}_{2}\bigr{)}\\ &={\hat{p}}_{2}^{2}+\gamma_{1}{\hat{q}}_{2}{\hat{p}}_{2}+\gamma_{2}\bigl{(}({\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}){\hat{p}}_{2}^{2}-\hat{\mathcal{C}}^{2}\bigr{)},\\[2.0pt] {\hat{I}}^{\prime}_{2}&={\hat{p}}_{1}^{2}+\gamma_{1}{\hat{q}}_{1}{\hat{p}}_{1}+\gamma_{2}\bigl{(}{\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}\bigr{)}{\hat{p}}_{1}^{2},\\ \end{split}$ (3.1)
@@
We observe that the corresponding classical functions $I_{2}$, $I_{2}^{\prime}$ and $\mathcal{K}_{i}$’s coming from the quantum operators (3.1) and (2.8) do not exactly coincide with $\mathcal{I}_{2}$, $\mathcal{I}_{2}^{\prime}$ and $\mathcal{L}_{i}$’s as obtained in [3]. Actually, they are related to these through $I_{2}=\mathcal{I}_{2}-\gamma_{2}\mathcal{C}^{2}$, $I_{2}^{\prime}=\mathcal{I}_{2}^{\prime}$, where $\mathcal{C}$ is given in (1.4). This explains why the classical counterpart of the quantum algebra (3.3) does not provide us with the algebra determined by the functions $\mathcal{L}_{i}$’s in [3]. These slight differences are also observable for higher values of $N$, and have their origin in the presence or not of the quantum angular momentum $\hat{\mathcal{C}}$.
We define the number and ladder operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ in (2.9) as
@@
The symmetry algebra of the quantum Hamiltonian ^$\hat{\mathcal{H}}_{2}$ becomes, over this basis, a cubic polynomial algebra $\mathfrak{sl}^{(3)}(2,\mathbb{R})$, which is just the so-called Higgs algebra [29] (see also [30, 31, 32, 33, 34] and references therein), as already pointed out in [5].
In order to construct a deformed Heisenberg algebra and compute the corresponding spectrum of ^$\hat{\mathcal{H}}_{2}$, we need to find the structure function $\Phi$ in (2.9), which turns out to be
@@ 3.1 Spectrum of the quadratic system
At this stage, we have all the information to compute the spectrum $E$ of ^$\hat{\mathcal{H}}_{2}$ in (2.4) algebraically. Using (3.5), it follows that the finite-dimensional representation of (2.12) yields the structure function $\Phi$ (2.13) as
@@
where we recall that $E$ and ${B}$ are the eigenvalues of ^$\hat{\mathcal{H}}_{2}$ and ${\hat{B}}$, respectively. Taking into account the two constraints in (2.14), the result can be summarized in the following
@@
The set of equations (2.14) yields four types of solutions for the representation-dependent constant $u$ and the spectrum of the Hamiltonian ^$\hat{\mathcal{H}}_{2}$ in (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$, that is, $u=u({n})$ and $E=E({n})$. Introducing these expressions into (3.6), the corresponding structure functions $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}=\{1,\dots,{n}\}$ are obtained. The final result is displayed in Table 1.
Table 1: The four types of solutions of the equations (2.14) with the corresponding representation dependent constant $u({n})$, spectrum $E({n})$ and structure function $\Phi(B,{n})$ (3.6) where ${n}\in\mathbb{N}^{\ast}$ and $B\in\{1,\dots,{n}\}$.
@@
Type III $u_{\rm III}=-\frac{1}{2}\left({n}-1+\frac{{\rm i}\gamma_{1}}{2\gamma_{2}}\right)$ $\ \displaystyle E_{\rm III}=-\frac{\gamma_{1}^{2}}{4\gamma_{2}}-\gamma_{2}({n}+1)^{2}$
@@
Type IV $u_{\rm IV}=-\frac{1}{2}\left({n}+1-\frac{{\rm i}\gamma_{1}}{2\gamma_{2}}\right)$ $\ \displaystyle E_{\rm IV}=-\frac{\gamma_{1}^{2}}{4\gamma_{2}}-\gamma_{2}({n}+1)^{2}$
@@
Some remarks are in order. The solutions of type III and IV give rise to the same spectrum, $E_{\rm III}=E_{\rm IV}$, although the structure functions $\Phi$ and hence the deformed oscillator algebras in (2.12) are different. The four types of spectra take real values whenever the parameter $\gamma_{1}$ is a pure imaginary number, keeping $\gamma_{2}$ as a real parameter. This suggests to introduce arbitrary real parameters $\alpha$ and $\beta$ as defined in (1.5), and appearing within the Zernike Hamiltonian (1.2). With these convention, we find that
@@
Different possibilities appear depending on the specific values of $\alpha$ and $\beta$. Quite unexpectedly, the quantum Zernike system corresponding to the values $\alpha=-1$ and $\beta=-2$ emerges as a very ‘special’ case, as the four types of solutions merge into the same, namely
@@
where ${n}\in\{1,2,\dots\}$ and ${B}\in\{1,\dots,{n}\}$. As $\Phi({B},n)>0$ holds, there is no restriction on the value of ${n}$, and the finite-dimensional representation is always unitary. Identifying the parameter ${n}$ with the principal quantum number $\mathfrak{n}$, we recover the expression
@@
given in [5] (cf. eq. (16)). Observe that, strictly speaking, ${n}\in\{1,2,\dots\}$, while $\mathfrak{n}\in\{0,1,2,\dots\}$, indicating that the difference relies on the ground state.
@@ 3.2 Application to the spherical and hyperbolic oscillators
In addition, it is worth remarking that, by means of a canonical transformation, it has been shown in [3] (see also [8]) that the classical Zernike system $\mathcal{H}_{\rm Zk}\equiv\mathcal{H}_{2}$ (2.3) can be regarded as an isotropic oscillator on the two-dimensional sphere $\mathbf{S}^{2}$ and the hyperbolic space $\mathbf{H}^{2}$. Within this interpretation, the imaginary coefficient $\gamma_{1}$ is related to the frequency $\omega$ in the form $\gamma_{1}=2{\rm i}\omega$, while the second one is just minus the constant curvature of the space $\gamma_{2}=-\kappa$. For the sake of completeness we recall the explicit form of $\mathcal{H}_{2}$ as a natural Hamiltonian in terms of a curved kinetic energy $\mathcal{T}_{\kappa}$ and a central potential $\mathcal{U}_{\kappa}\propto\tan^{2}(\sqrt{\kappa}\rho)/\kappa$:
@@
where $\rho$ is a geodesic radial coordinate, i.e., the distance between the particle and the origin on the space measured along the geodesic joining both points, while $\phi\in[0,2\pi)$ is an ordinary angle. Therefore, $\mathcal{H}_{2}$ covers the well-known two-dimensional isotropic spherical or Higgs oscillator on $\mathbf{S}^{2}$ ($\gamma_{2}<0$, $\kappa>0$), the hyperbolic system on $\mathbf{H}^{2}$ ($\gamma_{2}>0$, $\kappa<0$) and the usual one on the Euclidean plane $\mathbf{E}^{2}$ $(\gamma_{2}=\kappa=0)$ (see [35, 36, 37, 38, 39] and references therein). For the ‘standard’ values of the curvature $\kappa\in\{+1,0,-1\}$ we find that (3.9) gives
@@
From this perspective, the solutions of type III and IV in table 1 are not well-defined for the flat Euclidean case with $\kappa=\gamma_{2}\to 0$.
Although both quantum curved spherical and hyperbolic oscillators have already been fully solved in [35, 38] obtaining the corresponding spectrum and eigenfunctions, our algebraic procedure allows us to recover their spectrum straightforwardly, that shows deep differences between the spherical and hyperbolic cases, but both reproduce the Euclidean oscillator under the flat limit $\kappa\to 0$.
Let us focus on the solution of type I in table 1, set $\gamma_{1}=2{\rm i}$ i.e. $\beta=-2$ (this means to fix $\omega=1$) and express $\gamma_{2}=\alpha=-\kappa$, allowing any value for the curvature $\kappa$. We analyze separately the three particular systems according to $\kappa$ and for simplicity hereafter we drop the index ‘I’.
The case with $\kappa=0$ leads to the isotropic oscillator on $\mathbf{E}^{2}$ such that
@@
with ${n}\in\{1,2,\dots\}$, ${B}=\{1,\dots,{n}\}$ so that $\Phi>0$. Hence, as expected, the spectrum is linear in ${n}$, unbounded and with a constant difference $E(n+1)-E(n)=2$. This implies that there exists an infinite set of bound states.
The spherical or Higgs oscillator arises on $\mathbf{S}^{2}$ when $\kappa>0$ leading to the expressions
@@
where again ${n}\in\{1,2,\dots\}$, ${B}=\{1,\dots,{n}\}$ and $\Phi>0$. Now the spectrum is quadratic in ${n}$, with higher values than the Euclidean oscillator, and unbounded. Nevertheless, the difference of the energy between two consecutive states is no longer constant but increases linearly with ${n}$ as
@@
Consequently, there also exists an infinite number of bound states. We represent the first energies (3.11) for three values of the curvature and those of the Euclidean oscillator (3.10) in figure 1. Obviously, as the curvature grows the differences between the spherical and Euclidean spectra become larger.
@@
In contrast to (3.12), the difference of the energies is decreasing and turns into
@@
In order to relate these results to those previously deduced in [35, 38] and smoothly recover the Euclidean system, we require a positive energy spectrum $E(n)$ (3.13) and positive differences of two consecutive energies (3.14). The first condition leads to $2>|\kappa|n$, thus the lower value $n=1$ determines a limiting value for the curvature: $|\kappa|<2$. The second requirement implies that there exists a maximum value $n_{\rm max}$ for the positive integer $n\geq 1$, which is the greatest integer number fulfilling the constraint obtained by setting $n_{\rm max}=n+1\geq 2$ in (3.14), namely
$2\leq n_{\rm max}<\frac{1}{|\kappa|}+\frac{1}{2}.$
The single state (ground case) for the extreme value $n=n_{\rm max}=1$ can be included into a unique condition, such that both restrictions can be written as
$21\leq n_{\rm max}<\frac{1}{|\kappa|}+\frac{1}{2}\quad\Longrightarrow\quad|\kappa|<$ 2 . (3.15)
The implication is a direct consequence of taking the lowest $n_{\rm max}=1$. Moreover, observe that the specific values appearing in these relations depend not only on the curvature but also on the particular choice of the parameter $\beta=-2$ (hence on the frequency $\omega$). And, clearly, if $|\kappa|\to 0$, then $n_{\rm max}\to\infty$, recovering the Euclidean case. Provided that the value of $n_{\rm max}$ has been fixed, the relation (3.15) gives
$\begin{array}[]{ll}n_{\rm max}=1\!\!:&\quad\displaystyle\frac{2}{3}\leq|\kappa|<2.\\[10.0pt] n_{\rm max}>1\!\!:&\quad\displaystyle\frac{2}{2n_{\rm max}+1}\leq|\kappa|<\frac{2}{2n_{\rm max}-1},\end{array}$ (3.16)
that is, $32/5\leq|\kappa|<2/$ 3 for $n_{\rm max}=2$, $52/7\leq|\kappa|<2/$ 5 for $n_{\rm max}=3$, and so on. Thus one can look for a certain number of states choosing a value for $n_{\rm max}$ and obtain all possible compatible values for the curvature.
Let us now prove that the restriction (3.15) always ensure that $\Phi>0$ in (3.13) showing that $\bigr{(}1-|\kappa|(B-1)\bigl{)}>0$ and $\bigr{(}1-|\kappa|(n-B)\bigl{)}>0$. Taking into account that ${B}=\{1,\dots,{n}\}$ the maximum value for $B\geq 1$ is just $n_{\rm max}$ (3.15). For the first factor we find that
@@
Therefore and in agreement with [35, 38], the isotropic hyperbolic oscillator has a positive finite discrete spectrum with ${n}\in\{1,\dots,n_{\rm max}\}$ and there exist unbounded states which may be provided by one-dimensional hyperbolic Pöschl–Teller potentials, as pointed out in Theorem 6 in [38]. We illustrate these results in figure 2, plotting the discrete spectra for three values of $n_{\rm max}$ (so choosing a value for the curvature) which are always lower than the Euclidean ones (3.10).
[…]

### v2

@@ 2 Generalized quantum Zernike Hamiltonians
[…]
@@
[…]
@@
  • Finally, using the eigenvalues $E$ of $\hat{\mathcal{H}}_{N}$ and B{B} of $^{\hat{B}}$, we consider a finite-dimensional representation of (2.12), that is,
@@
    for a natural number ${n}\in\{1,2,\dots\}$. This set of algebraic equations provides us with constraints on the spectrum and the representation-dependent constant $u$. Hence the solutions take the form u=u(n)u=u({n}) and E=E(n)E=E({n}), thus leading to $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}\in\{1,\dots,{n}\}$. If $\Phi(B,{n})>0$, the finite-dimensional representation is also unitary. The parameter n{n}, which characterizes the dimension of the subspace, i.e. of dimension n+1{n}+1, determines the principal quantum number $\mathfrak{n}$.
In the following sections 3 and 4, we work out the quadratic and cubic generalized quantum Zernike Hamiltonians $\hat{\mathcal{H}}_{N}$ (2.3), corresponding to the values $N=2$ and $N=3$, respectively. For each case, we obtain the quantum symmetries (2.7), the polynomial algebras (2.8) and (2.9), together with the deformed oscillator algebra (2.12). From the latter, the solutions for the possible spectra determined by the equations (2.14) are deduced and analyzed. Furthermore, such new results for $N=3$ are also interpreted as superintegrable perturbations of the isotropic oscillator on the two-dimensional sphere, hyperbolic and Euclidean spaces; recall that the usual isotropic Euclidean oscillator is just the case with $N=1$, thus with a single parameter $\gamma_{1}$.
In addition, following a similar construction, the fourth-order with $N=4$ and the fifth-order with $N=5$ systems are explicitly solved. To facilitate the reading of the paper, the corresponding results are presented in Appendices A and B, respectively. While we have not yet fully derived the general expressions for the deformed oscillator algebra (2.12) along with the corresponding spectra for arbitrary $N$, we present two conjectures in section 5 that address these problems. These conjectures serve as a foundation for future work to establish their proofs.
@@ 3 The proper (quadratic) quantum Zernike Hamiltonian
Let us consider the quadratic case of $\hat{\mathcal{H}}_{N}$ (2.3) with $N=2$ and arbitrary parameters $\gamma_{1}$ and $\gamma_{2}$. The second-order in the momenta quantum symmetries (2.6) commuting with $\hat{\mathcal{H}}_{2}$ (2.7) are found to be
[…]
@@
We observe that the corresponding classical functions $I_{2}$, $I_{2}^{\prime}$ and $\mathcal{K}_{i}$’s coming from the quantum operators (3.1) and (2.8) do not exactly coincide with $\mathcal{I}_{2}$, $\mathcal{I}_{2}^{\prime}$ and $\mathcal{L}_{i}$’s as obtained in [4]. Actually, they are related to these through $I_{2}=\mathcal{I}_{2}-\gamma_{2}\mathcal{C}^{2}$, $I_{2}^{\prime}=\mathcal{I}_{2}^{\prime}$, where $\mathcal{C}$, $\mathcal{I}_{2}\equiv\mathcal{I}$ and $\mathcal{I}^{\prime}_{2}\equiv\mathcal{I}^{\prime}$ are given in (1.3), (1.4) and (1.5), respectively and, therefore, they are constants of the motion of the classical Hamiltonian $\mathcal{H}_{2}$ (1.8). This explains why the classical counterpart of the quantum algebra (3.3) does not provide us with the Poisson symmetry algebra determined by the functions $\mathcal{L}_{i}$’s in [4], given by formulae (1.6) and (1.7). These slight differences are also observable for higher values of $N$, and have their origin in the presence or absence of the quantum angular momentum operator $\hat{\mathcal{C}}$ in the quantum symmetries.
[…]
@@
The symmetry algebra of the quantum Hamiltonian $\hat{\mathcal{H}}_{2}$ becomes, over this basis, a cubic polynomial algebra $\mathfrak{sl}^{(3)}(2,\mathbb{R})$, which is just the so-called Higgs algebra [17] (see also [34, 35, 36, 37, 38] and references therein), as already pointed out in [6].
In order to construct a deformed Heisenberg algebra and compute the corresponding spectrum of $\hat{\mathcal{H}}_{2}$, we need to find the structure function $\Phi$ in (2.9), which turns out to be
@@ 3.1 Spectrum of the quadratic system
At this stage, we have all the required information to compute the spectrum $E$ of $\hat{\mathcal{H}}_{2}$ in (2.4) algebraically. Using (3.5), it follows that the finite-dimensional representation of (2.12) yields the structure function $\Phi$ (2.13) as
@@
where we recall that $E$ and B{B} are the eigenvalues of $\hat{\mathcal{H}}_{2}$ and $^{\hat{B}}$, respectively. Taking into account the two constraints in (2.14), the result can be summarized in the following
@@
The set of equations (2.14) yields four types of solutions for the representation-dependent constant $u$ and the spectrum of the Hamiltonian $\hat{\mathcal{H}}_{2}$ in (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$, that is, u=u(n)u=u({n}) and E=E(n)E=E({n}). Introducing these expressions into (3.6), the corresponding structure functions $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}=\{1,\dots,{n}\}$ are obtained. The final result is displayed in Table 1.
Table 1: The four types of solutions of the equations (2.14) with the corresponding representation dependent constant u(n)u({n}), spectrum E(n)E({n}) and structure function $\Phi(B,{n})$ (3.6) where ${n}\in\{1,2,\dots\}$ and $B\in\{1,\dots,{n}\}$.
@@
[…]
@@
[…]
@@
Some remarks are in order. It can be directly verified that the four types of solutions in table 1 satisfy both equations (2.14), so that the factors $\Phi_{1}$ and $\Phi_{2}$ in (3.6) cancel out as follows:
$\begin{array}[]{ll}\Phi_{1}(0,E_{\rm I},u_{\rm I})=\Phi_{2}(n+1,E_{\rm I},u_{\rm I})=0,\quad&\Phi_{1}(n+1,E_{\rm II},u_{\rm II})=\Phi_{2}(0,E_{\rm II},u_{\rm II})=0,\\[4.0pt] \Phi_{2}(0,E_{\rm III},u_{\rm III})=\Phi_{2}(n+1,E_{\rm III},u_{\rm III})=0,\quad&\Phi_{1}(0,E_{\rm IV},u_{\rm IV})=\Phi_{1}(n+1,E_{\rm IV},u_{\rm IV})=0.\end{array}$
The solutions of type III and IV give rise to the same spectrum, $E_{\rm III}=E_{\rm IV}$, although the structure functions $\Phi$ and hence the deformed oscillator algebras in (2.12) are different. The four types of spectra take real values whenever the parameter $\gamma_{1}$ is a pure imaginary number, keeping $\gamma_{2}$ as a real parameter. This suggests to introduce arbitrary real parameters $\alpha$ and $\beta$ as defined in (1.9), and appearing within the Zernike Hamiltonian (2.2). With this convention, we find that
@@
Different possibilities appear depending on the specific values of $\alpha$ and $\beta$. In this regard, it should be noted that if the value $\alpha=\gamma_{2}=0$ is allowed, the spectrum $E_{\rm III}=E_{\rm IV}$ is not well-defined. This fact is quite important, since this case corresponds to the system with $N=1$, which actually corresponds to the isotropic oscillator on the Euclidean plane, that we study next. By contrast, the two spectra $E_{\rm I}$ and $E_{\rm II}$ do cover the system with $\alpha=\gamma_{2}=0$. Moreover, we also point out that $E_{\rm II}$ can be related to $E_{\rm I}$ by applying a ‘shift’ $n\to(n+2)$ along with the change $\beta\to-\beta$ to $E_{\rm I}$.
Furthermore, and quite unexpectedly, the quantum Zernike system corresponding to the values $\alpha=-1$ and $\beta=-2$ arises as a very ‘special’ case, as the four types of solutions merge into the same, namely
@@
where ${n}\in\{1,2,\dots\}$ and ${B}\in\{1,\dots,{n}\}$. As $\Phi({B},n)>0$ holds, there is no restriction on the value of n{n}, and the finite-dimensional representation is always unitary. Identifying the parameter n{n} with the principal quantum number $\mathfrak{n}$, we recover the expression
@@
given in [6] (cf. eq. (16)). Observe that, strictly speaking, ${n}\in\{1,2,\dots\}$, while $\mathfrak{n}\in\{0,1,2,\dots\}$, indicating that the difference relies merely on the ground state.
@@ 3.2 Application to the spherical and hyperbolic oscillators
In addition, it is worth remarking that, by means of a canonical transformation, it has been shown in [4] (see also [9]) that the classical Zernike system $\mathcal{H}_{\rm Zk}\equiv\mathcal{H}_{2}$ (2.3) can be regarded as an isotropic oscillator on the two-dimensional sphere $\mathbf{S}^{2}$ and the hyperbolic space $\mathbf{H}^{2}$. Within this interpretation, the imaginary coefficient $\gamma_{1}$ is related to the frequency $\omega$ in the form $\gamma_{1}=2{\rm i}\omega$, while the second one is just minus the constant curvature of the space $\gamma_{2}=-\kappa$. For the sake of completeness, we recall the explicit form of $\mathcal{H}_{2}$ as a natural Hamiltonian in terms of a curved kinetic energy $\mathcal{T}_{\kappa}$ and a central potential $\mathcal{U}_{\kappa}\propto\tan^{2}(\sqrt{\kappa}\rho)/\kappa$:
@@
where $\rho$ is a geodesic radial coordinate, i.e., the distance between the particle and the origin on the space measured along the geodesic joining both points, while $\phi\in[0,2\pi)$ is an ordinary angle. Therefore, $\mathcal{H}_{2}$ covers the well-known two-dimensional isotropic spherical or Higgs oscillator on $\mathbf{S}^{2}$ ($\gamma_{2}<0$, $\kappa>0$), the hyperbolic system on $\mathbf{H}^{2}$ ($\gamma_{2}>0$, $\kappa<0$) and the usual one on the Euclidean plane $\mathbf{E}^{2}$ $(\gamma_{2}=-\kappa=0)$ (see [39, 40, 41, 42, 43] and references therein). For the ‘standard’ values of the curvature $\kappa\in\{+1,0,-1\}$ we find that (3.9) gives
@@
From this perspective, the solutions of type III and IV in Table 1 are not well-defined for the flat Euclidean case with $\kappa=-\gamma_{2}\to 0$, as we have already mentioned.
Although both quantum curved spherical and hyperbolic oscillators have already been fully solved in [39, 42], obtaining the corresponding spectrum and eigenfunctions, our algebraic procedure allows us to recover their spectrum in straightforward manner, that shows deep differences between the spherical and hyperbolic cases. Nevertheless, both reproduce the Euclidean oscillator under the flat limit $\kappa\to 0$, which is only allowed for types I and II in Table 1.
Let us focus on the solution of type I in Table 1, setting $\gamma_{1}=2{\rm i}$ i.e. $\beta=-2$ (this means to fix $\omega=1$) and express $\gamma_{2}=\alpha=-\kappa$, allowing any value for the curvature $\kappa$. We analyze separately the three particular systems according to $\kappa$ and, for simplicity, hereafter we drop the index ‘I’.
The case with $\kappa=0$ leads to the isotropic oscillator on $\mathbf{E}^{2}$ (hence, the case with $N=1$ in (2.3)), such that
@@
with ${n}\in\{1,2,\dots\}$, ${B}=\{1,\dots,{n}\}$ so that $\Phi>0$. Thus, as expected, the spectrum is linear in n{n}, unbounded and with a constant difference E(n+1)E(n)=2E(n+1)-E(n)=2. This implies that there exists an infinite set of bound states.
The spherical or Higgs oscillator arises on $\mathbf{S}^{2}$ when $\kappa>0$, leading to the expressions
@@
where, again, ${n}\in\{1,2,\dots\}$, ${B}=\{1,\dots,{n}\}$ and $\Phi>0$. Now the spectrum is quadratic in n{n}, with higher values than the Euclidean oscillator, and unbounded. Nevertheless, the difference of the energy between two consecutive states is no longer constant but increases linearly with n{n} as
@@
We conclude that, in this case, there also exists an infinite number of bound states. We represent the first energies (3.11) for three values of the curvature, and those of the Euclidean oscillator (3.10) in figure 1. As the curvature grows, the differences between the spherical and Euclidean spectra become larger.
@@
In contrast to (3.12), the difference of the energies is decreasing and becomes
@@
In order to relate these results to those previously deduced in [39, 42] and recover the Euclidean system by a limiting process, we require a positive energy spectrum $E(n)$ (3.13) and positive differences of two consecutive energies (3.14). The first condition leads to $2>|\kappa|n$, thus the lower value $n=1$ determines a limiting value for the curvature: $|\kappa|<2$. The second requirement implies that there exists a maximum value $n_{\rm max}$ for the positive integer $n\geq 1$, which is the greatest integer number ensuring that the relation (3.14) is always positive, namely
$E(n_{\rm max})-E(n_{\rm max}-1)=2-|\kappa|\big(2(n_{\rm max}-1)+1\big)>0.$
This constraint gives rise to
$1\leq n_{\rm max}<\frac{1}{|\kappa|}+\frac{1}{2}.$ (3.15)
Notice that the single state (ground case) for the extreme value $n_{\rm max}=1$ is also included, since it consistently implies that $|\kappa|<2$. Moreover, observe that the specific values appearing in relations (3.13)–(3.15) depend not only on the (negative) curvature, but also on the particular choice of the parameter $\beta=-2$ (and, therefore, on the frequency $\omega$). And if $|\kappa|\to 0$, then $n_{\rm max}\to\infty$, recovering the Euclidean case.
Let us now prove that the restriction (3.15) always ensures that $\Phi>0$ in (3.13), showing that $\bigr(1-|\kappa|(B-1)\bigl)>0$ and $\bigr(1-|\kappa|(n-B)\bigl)>0$. Taking into account that ${B}=\{1,\dots,{n}\}$, the maximum value for $B\geq 1$ is just $n_{\rm max}$ (3.15). For the first factor we find that
@@
Figure 1: The discrete spectrum (3.11) of the spherical oscillator on $\mathbf{S}^{2}$ for the first ten states $1\leq n\leq 10$ according to three values of the curvature $\kappa\in\{0.5,\,0.25,\,0.15\}$ starting from the upper dots together with the Euclidean spectrum (3.10) with $\kappa=0$ in the lowest values.
Figure 2: The finite discrete spectrum (3.13) of the hyperbolic oscillator on $\mathbf{H}^{2}$ for the three values of the curvature $\kappa\in\{-0.25,\,-0.16,\,-0.12\}$ corresponding to $n_{\rm max}\in\{4,\,6,\,8\}$ (3.15) starting from the lower dots, respectively, together with the Euclidean spectrum (3.10) with $\kappa=0$ in the highest values.
Therefore, in agreement with [39, 42], the isotropic hyperbolic oscillator has a positive finite discrete spectrum with ${n}\in\{1,\dots,n_{\rm max}\}$, and there exist unbounded states which may be provided by one-dimensional hyperbolic Pöschl–Teller potentials, as pointed out in Theorem 6 in [42]. We illustrate these results in figure 2, plotting the discrete spectra for three values of $n_{\rm max}$ (so choosing a value for the curvature) which are always lower than the Euclidean ones (3.10).